OpenStudy (bananas1234):

What is -squrt90 expressed in simplified form?

OpenStudy (bananas1234):

\[-\sqrt{90}\]

OpenStudy (bananas1234):

@zepdrix

OpenStudy (bananas1234):

@mathmale

OpenStudy (bananas1234):

@Directrix

OpenStudy (bananas1234):

@jagr2713

jhonyy9 (jhonyy9):

rewrite the 90 = 9*10 = 3^2 *10 in this way you get -sqrt90 = -sqrt(10*3^2) = -3sqrt10 hope helped

OpenStudy (3mar):

As @jhonyy9 says, try to expand 90 under the square root...